OpenStudy (anonymous):

Simplify 5 square root of 40

OpenStudy (anonymous):

You're starting with this: \[5 \sqrt{40}\]To simplify you need to find a factor of 40 that is a perfect square. You can then pull the factor out, square root it, and put it in front. So: \[5 \sqrt{40} = 5 \sqrt{4 \times 10} = 5 \times 2 \sqrt{10} = 10 \sqrt{10}\]